OpenStudy (anonymous):

Decide if the statement is always, sometimes, or never true. A parabola has 2 x-intercepts. Answer always sometimes never

OpenStudy (anonymous):

Sometimes — y=x^2 has one, and y=x^2+1 has none at all.

OpenStudy (anonymous):

But, say y=x^2-1 does have two.

OpenStudy (anonymous):

@mitodoteira I was going to say that